Last year, X Company sold 68,000 units of its only product for $18.00 each. Total costs were as follows:

Cost of goods sold

Variable $407,320

Fixed 127,160

Selling and administrative

Variable $96,560

Fixed 101,320

At the end of the year, a company offered to buy 4,710 units of the product but only for $12.00 each. X Company had the capacity to produce the additional units, and even though there would have been no additional selling and administrative costs, it rejected the offer.

5. If X Company had accepted the special order, firm profits would have increased by

6. X Company's production manager thought that the special order units might have required additional direct material and direct labor costs per unit of $0.84 and $0.28, respectively, and the rental of special equipment for $1,500. The combined effect of these changes would have been to reduce the special order profit by

7. X Company's marketing manager felt that in order for regular sales to continue at 68,000 units this year, the company would have had to reduce the regular selling price this year to $17.65. The result of this price reduction would have been to reduce X Company's profits this year by
Answer #1

Solution 5:

Variable cost of goods sold per unit = $407,320 / 68000 = $5.99 per unit

If X company accept the special order then profit will increase by = ($12 - $5.99) * 4710 = $28,307

Solution 6:

Reduction in special order profit by = ($0.84 + $0.28) * 4710 + $1,500 = $6,775

Solution 7:

Price reduction would reduce X company profit by = (68000*$0.35) = $23,800